You are rolling a pair of dice, one red and one green. What is the probability of the following outcomes:
a. The sum of the two numbers you roll from the dice is 11.
b. The sum of the two numbers you roll is 7.
c. The sum of the two numbers you roll is 13.


Answer by stanbon(48551)

You are rolling a pair of dice, one red and one green. What is the probability of the following outcomes:
-----------
You should write out the sample space as a 6 by 6 array showing the
distribution of sums of the two die.
----------------------------
a. The sum of the two numbers you roll from the dice is 11.
2/36 = 1/18
-------------------
b. The sum of the two numbers you roll is 7.
6/36= 1/6
-------------------
c. The sum of the two numbers you roll is 13.
0/36 = 0
=============================
Cheers,
Stan H.